OpenStudy (anonymous):

Can someone check my work? 1. Convert 7pi/6 radians to degrees. My answer: 210

OpenStudy (anonymous):

@ayeshaafzal221

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

\[\frac{ 7 \pi }{ 6 } \times \frac{ 180 }{ \pi } =210\]